OpenStudy (loser66):

Let \(X_n\) be a Cauchy sequence. Suppose that for every \(\varepsilon >0\), there is some \(n>1/\varepsilon\) such that \(|X_n|<\varepsilon\). Prove that \(X_n\rightarrow 0\) Please, help.

OpenStudy (loser66):

My attempt: Since \(X_n\) is a Cauchy sequence, Let \(\varepsilon >0\), then for m >n, there exists \(N_1 >0\) such that \(\forall~~m, n>N_1\) \(|X_n-X_m|< \varepsilon/2\) And we have \(|X_n|= |X_n -X_k+X_k|\leq |X_n-X_k|+|X_k|\) By above, if k > n, then the first term from the right hand side \(<\varepsilon/2\) All we need is manipulate \(|X_k| <\varepsilon /2\). I stuck here. I don't know how to argue.

OpenStudy (loser66):

I know I need play around with the condition \(n > 1/\varepsilon\) such that \(|X_n|<\varepsilon\) to get what I want, but I don't know how. Someone helps me, please

OpenStudy (anonymous):

Why not just choose \(\varepsilon\) small enough so that \(\frac{1}{\varepsilon} > N_1\)?

OpenStudy (anonymous):

Suppose you chose \(\varepsilon\), then using the condition choose: \(\displaystyle\varepsilon'= \min\ \left\{\frac{1}{N_1}, \frac{\varepsilon}{2}\right\}\) Then you have \(\displaystyle|X_n - X_k| + |X_k| < \frac{\varepsilon}{2} + \varepsilon' < \varepsilon\)

OpenStudy (loser66):

Got it, thanks a lot. :)